If a charge of magnitude 7.5 x 10-17 C is being held in place 2 x 10-8 m

from a charge of -9.0 x 10-19C which is also being held in place, what is the electric

potential energy of the system?



1
Expert's answer
2021-04-05T11:11:25-0400

We can find the electric potential energy of the system as follows:


"U=\dfrac{kQ_1Q_2}{r},""U=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot7.5\cdot10^{-17}\ C\cdot(-9\cdot10^{-19}\ C)}{2\cdot10^{-8}\ m},""U=-3.04\cdot10^{-17}\ J."